# The Fast and the Frame-Dragging: Efficient waveforms for asymmetric-mass eccentric equatorial inspirals into rapidly-spinning black holes

### v1

@@ III.3 Mode selection and waveform summation
While the inclusion of spin on the primary introduces additional complexity to the generation of inspiral trajectories and mode amplitudes, once these have been addressed, the waveform may be constructed in much the same way as for spinless systems. We review this procedure below, which is largely unchanged except for the incorporation of the continuous ODE solution described in Section III.1.3; see Katz:2021yft and Speri:2023jte for more information regarding these stages of waveform generation in the time domain and frequency domain, respectively.
@@
Before performing the summation in Eq. 1, we apply a mode selection process to reduce the number of modes that must be summed without sacrificing accuracy, which proceeds as follows. First, at each sparse trajectory point, we compute the mode power $|A_{\ell mn}|^{2}$ for each waveform mode in our amplitude data grids. We then sort the list of mode power values at each individual point, perform a cumulative summation over this list and truncate the sum once it exceeds a fraction $(1-\kappa)$ of the total power. We assume that modes that do not pass this cut-off do not strongly contribute to the overall waveform. However, to ensure that the waveform is smooth with respect to time, we take the union of the sets of modes that pass the individual cut-offs and retain only these modes for waveform summation. The tuning parameter $\kappa$ controls the accuracy of the waveform with respect to this procedure; in line with previous work, we set $\kappa=10^{-5}$ by default. We have verified that this value is a sensible choice for a variety of EMRIs and IMRIs, as discussed in Section V.6.
@@
After mode selection, we perform the summation in Eq. 1 over the remaining sets of mode indices. For time-domain waveforms, we interpolate the sparse quantities $\mathcal{A}_{\ell mn}$ and $\Phi_{mn}$ (with cubic splines and the continuous ODE solution respectively) on a uniform grid of time values of spacing $\mathrm{d}t$, summing the interpolated waveform modes together according to Eq. 1. Frequency-domain waveforms are obtained via the approach described in Speri:2023jte , making use of the stationary phase approximation. As this approach assumes that the frequency evolution of each mode is described by cubic splines (such that the stationary phase relation has analytic roots), we retain this structure rather than incorporating the continuous solution depicted in Fig. 2. The frequency-domain waveform also requires higher-order frequency derivatives to approximate the Fourier phase in each bin; as the derivatives of this cubic spline are less accurate than those obtained with the continuous solution, we rewrite this part of the frequency-domain summation to use the coefficients output by the ODE solver. We found that this significantly improved the consistency of the time-domain and frequency-domain waveforms. Modifying the root-finding procedure of the stationary phase approximation to incorporate our more accurate solution for the frequency evolution (e.g. by taking the analytic root as an input to Newton’s method, which will converge in 1-2 iterations) will further improve this consistency at minimal computational cost, and is left to future work.
@@ IV Validation of model accuracy
With our waveform generation framework fully specified, we will now explore the accuracy of our waveform model at adiabatic order. This is made difficult by lack of accurate waveform templates to compare against: numerical relativity simulations have not been computed (however, see recent efforts in Refs. Lousto:2022hoq ; Wittek:2024gxn ; Wittek:2024pis ) in this region of the parameter space (especially for the number of orbital cycles [$\sim\mathcal{O}(\nu^{-1})$] required for useful comparisons to be made), and PN models are not trustworthy in the strong-field region of our domain where confirming the accuracy of our model is most pertinent. We therefore opt for a different approach in which we quantify significant sources of systematic error in our model (which in combination are responsible for any differences between our model and “exact” adiabatic EMRI waveforms), with the aim of understanding how these errors propagate to the accuracy of our constructed waveforms. In this way, we can roughly quantify the faithfulness of our model with respect to an “error-free” adiabatic model without the need for independent waveforms to compare against. Throughout this section, we will examine each component of our waveform generation in turn: we first consider the trajectory module (Section IV.1), followed by the amplitude module (Section IV.2) and conclude with an examination of how these systematics impact accuracy at the waveform template level (Section IV.3).
In addition to this self-contained examination of our model, we also verify that it correctly matches other models in limiting cases in Appendix E. In combination with the systematic validation that we perform throughout this section, the results of these comparisons confirm that our model produces accurate waveforms for spinning and/or equatorial IMRIs and EMRIs across its domain of validity of $|a|\leq 0.999$ and $e_{0}\leq 0.9$. In this process, we identify that our model is least accurate for: $a>0.99$ and $p-p_{\mathrm{sep}}\lesssim 1$; and $e\gtrsim 0.85$. These are regions where self-force data vary rapidly and are most challenging to interpolate.
@@ IV.1 Trajectory validation
In this subsection, we verify that our model for the inspiral trajectory of the CO accurately traces the orbital phasing of the system to adiabatic order over the long timescales typical of the asymmetric-mass binary systems observable by LISA. As our model only incorporates adiabatic contributions to the inspiral evolution, one could argue that we need only construct the phase evolution to an accuracy of $\mathcal{O}(1)$ radian, as we cannot hope to match trajectories that include post-adiabatic contributions to better than this scale (and typically, ignoring post-adiabatic terms regularly leads to $10$-$100$ radians of phase error Isoyama:2021jjd ; Wardell:2021fyy ; Burke:2023lno ; Lynch:2024ohd , at least for the quasi-circular EMRIs). However, we can assume that post-adiabatic effects will be implemented in our framework once they are readily obtainable (as they are necessary to achieve sufficient accuracy for data analysis purposes), so this is not a useful requirement for us to satisfy. In this work, our focus remains on the accuracy of our model for the adiabatic contribution to the inspiral. Quantitatively assessing how phase errors translate into systematic biases in EMRI parameter estimation is not straightforward. However, in Ref. Burke:2023lno , which examined quasi-circular inspirals into non-spinning MBHs, a phase error of $\sim 10^{-1}$ radians was sufficiently small for any resulting biases to remain subdominant to statistical uncertainties. Our aim therefore is to ensure that our trajectory model reproduces inspiral phasing with at least this level of accuracy.
@@
Figure 5: Relative error in $\hat{f}_{p}^{(0)}$ between an interpolation of our data grid and the corresponding data of Ref. Fujita:2020zxe , available from the BHPC data lake BHPClub . The two datasets agree to better than $8$ digits over the majority of the parameter space. Here we examine prograde orbits with $a=0.5$; we observe similar agreement for other spin values $a\in[-0.9,0.9]$.
@@
The first of these datasets was computed by the authors of Ref. Fujita:2020zxe (built on the BHPC code developed in Refs. Fujita:2004rb ; Fujita:2005kng ; Fujita:2009us ) and is publicly available on the BHPC website BHPClub . It broadly and densely covers a large fraction of our parameter space. The authors accompany their data with an error estimate informed by their truncation of mode-summations. For $p\lesssim p_{\mathrm{sep}}+1$, the estimated error becomes prohibitively large for validation purposes — in this region, the discrepancy between our data and theirs grows proportionally to the predicted error. We therefore excise this region of the parameter space from this comparison (and will examine it separately below with a separate dataset). The region of the BHPC data we retain extends to $p\sim 25$ and $e=0.925$, providing good coverage of a large fraction of our grid. Beyond $p\sim 25$, we validate our data separately with BHPC’s PN results (see Section E.2). The relative error between an interpolation of our data and the BHPC data for $\hat{f}_{p}^{(0)}$ is shown in Fig. 5 — for brevity, we do not include the comparison with $\hat{f}_{e}^{(0)}$, which follows similar behaviour. We attain excellent agreement with this dataset across the parameter space, with relative error remaining below $\sim 10^{-8}$ over the vast majority of the grid. Above $e\sim 0.8$, errors steadily rise to $\sim 10^{-4}$. This is expected: mode-sums with respect to $n$ converge slowly in this region, requiring one to compute and sum over hundreds of $n$-modes. The large $n$-harmonics become increasingly more difficult to resolve at floating point precision and introduce their own source of error if they are under-resolved (similar to how it is incredibly difficult to calculate the 100th harmonic in a Fourier series expansion). For a single $n$-mode, this error is typically subdominant to the total error, but if the mode-sum is not truncated early enough, the accumulation of these small errors can saturate the overall relative accuracy of the results. As both our data and that of the BHPC are impacted by this limitation, it is not clear which of the two have been computed more accurately in this high eccentricity region, only that the two datasets are consistent to $\sim 4$ digits of precision. This comparison is performed for prograde orbits with $a=0.5$ — we find similarly good agreement across all spin values $(a=0.1,0.3,0.5,0.7,0.9)$ for both prograde and retrograde configurations, confirming that we have accurately computed and interpolated our $\hat{f}_{p,e}^{(0)}$ data grids over this part of the parameter space.
Figure 6: Relative error in $\hat{f}_{p}^{(0)}$ between an interpolation of our data grid and a dataset computed with the GREMLIN code. The four panels shown correspond to four slices in eccentricity that follow the tapering scheme described in Appendix B. The two datasets generally agree to6similar-toabsent6$\sim$ 6∼ 6 digits for $p-p_{\mathrm{sep}}\gtrsim 10^{-2}$, increasing to $\sim 4$ digits below this value. Some data quality issues are also evident in the lower right panel; see main text for discussion.
@@
To inspect the accuracy of our data grids in the strong-field region $p\rightarrow p_{\mathrm{sep}}$, we also compare against data computed with the GREMLIN code, independently developed by author Hughes and previous collaborators. (See Ref. Drasco:2005kz for an overview of this code’s foundations, and Refs. ThroweThesis ; OSullivan:2014ywd for discussion of updates to notation and methods that this code uses. Additional recent optimizations and algorithmic improvements, including a port to the Open Science Grid osg07 , will be described in a future publication, accompanied by a public release after removing proprietary code currently used by this package.) This data is more densely sampled in this region than our grids in $a$ and $p$; comparisons are done on $4$ slices in eccentricity that follow a taper similar to the one we include in our grid reparametrization. We distinguish forcing functions computed with few and GREMLIN with superscript “FEW” and “SAH” respectively. The relative error between an interpolation of our data and each of these slices is shown in Fig. 6 for the case of $\hat{f}_{p}^{(0)}$. We observe excellent agreement (with relative errors less than one part in a million) for $p-p_{\mathrm{sep}}\gtrsim 10^{-2}$. Closer to the separatrix, we observe a larger discrepancy between the datasets, with relative errors reaching $\sim 10^{-3}$; this is expected as (especially at larger eccentricities) it is computationally challenging to compute $\hat{f}_{p,e}^{(0)}$ in this region, and our interpolation error begins to dominate at high values of $a$ (as indicated by the alternating bands of larger error with respect to $a$), as we will examine further below. However, EMRIs typically complete relatively few orbital cycles in this region of the parameter space (compared to their total number of in-band cycles) before crossing the separatrix, so even relatively large errors will not significantly impact the phase accuracy of the trajectory.
@@
While the deviations observed are due in part to errors in the interpolation of our data grids, they are also impacted by differences in the convergence criteria associated with each software. The gremlin computations truncate at either a relative precision of $10^{-7}$ or up to $n=250$ closest to the separatrix, whereas pybhpt cuts off at either a relative precision of $10^{-8}$ or $n=300$; as these higher $n$-modes can contain significant errors due to limited numerical precision, truncating the mode-sum at different values of $n$ may yield results that differ substantially beyond the first few digits. Moreover, relative precision cut-offs may not be directly comparable between datasets, as this quantity may be estimated in different numerical frameworks. Likewise, the BHPC data were computed with mode-sums truncated at up to $n=1000$ or $\ell=25$ for the (highly-eccentric) strongest-field orbits to ensure a relative precision at a level of $10^{-6}$ Fujita:2020zxe , which may again yield significant differences with our data (such as those observed for the highest eccentricities in Fig. 5). Errors due to these differences in methodology as a function of parameter space are not well understood at present; further investigation of these systematics is beyond the scope of this work khalvati2025sys (but is worthy of future study, particularly for eccentric and inclined orbits).
@@
Two other significant features are present in the heatmaps shown in Fig. 6. The first is that for $a\rightarrow 0.999$, we observe a significant increase in relative error in alternating bands of constant $a$. These bands correspond to $a$-values lying on grid nodes (low interpolation error) and exactly between grid nodes (high interpolation error) respectively, and their existence suggests that our grid density with respect to $a$ is insufficient near $p_{\mathrm{sep}}$ (even despite our reparameterization, which concentrates grid points in this region). As we will see later in this subsection, this has little impact on the accuracy of our trajectories. The second feature is observed for the largest eccentricity slice (bottom-right panel of Fig. 6), $a\in[0,0.5]$ and $p-p_{\mathrm{sep}}\lesssim 1$, where a patch of larger relative error is observed. These points correspond to $e\sim 0.5$ near the separatrix (due to the eccentricity taper with respect to $e_{\mathrm{max}}=0.9$, see Appendix B) and the larger errors are a consequence of ill-converged mode-sums typical of more eccentric orbits near $p=p_{\mathrm{sep}}$. As this eccentricity slice is at the edge of our grid, trajectories typically do not perform many orbital cycles in this region before their eccentricities decay to lower values where this numerical precision issue is absent; we therefore do not anticipate this will significant impact the accuracy of our model. In this eccentricity slice, a steadily-growing relative error is also observed for $a\rightarrow-0.999$, which corresponds to $e\rightarrow 0.9$ and is explained by the differences in convergence criteria between implementations (as described above).
@@
Taking these comparisons with independent datasets as a reference point, we assess the impact of interpolation error on our estimation of $f_{p,e}^{(0)}$ as a function of the inspiral parameters. As our data grids have dimensions of the form $12^{n}+$ 1 for some positive integer $n$, we can construct a grid of dimension $12^{n-1}+$ 1 (halving the point density) by omitting every second point in the grid, which we will refer to as “down-sampling” the grid by a factor of $2$. We can then evaluate this down-sampled grid at the omitted points, placing the upper bound on the interpolation error for the full grid across the parameter space. The comparison is performed in Fig. 7 for $a\sim 0.998$, the spin value at which we find the interpolation errors to be most significant (in line with what is observed in Fig. 6); again, we omit our comparison with $\hat{f}_{e}^{(0)}$ for brevity, which yields similar error estimates. Over the majority of the parameter space, this interpolation error estimate exceeds those shown in Figs. 6 and 5 where we compare to independent datasets. This means that for a trajectory model built with data grids down-sampled by a factor of 2 (or higher), interpolation error due to insufficient point density will be the dominant contribution to orbital dephasing.
@@
We will now use this result to roughly quantify the accuracy of our model. We consider models built from data-grids down-sampled by factors of $2$, $4$ and $8$, and compare phase trajectories constructed with these down-sampled models to those produced by our full-scale model (which has the full grid density). In the top panel of Fig. 8, we plot histograms for dephasings between each down-sampling scenario and our full-resolution model for a random sample of $10^{3}$ inspirals with $p_{0}$ tuned such that they inspiral after 4 years (see Appendix C for details of the random-draw procedure). We observe that these dephasings display power-law convergence (i.e., in line with the power of $2$ of the down-sampling factor) as a function of grid density. By fitting this convergence trend between grid density and dephasing for each data point, we then extrapolate to obtain an estimate of the phase error at the full grid density, resulting in the dashed histogram in the top panel of Fig. 8. Taking this extrapolation as a rough estimate of our model’s dephasing with respect to a “perfect” adiabatic inspiral model, we predict that our model attains a phase accuracy of $\sim 10^{-4}\mathrm{\leavevmode\nobreak\ rad}$ over the majority of the parameter space, with a higher tail extending to $\sim 10^{-2}\mathrm{\leavevmode\nobreak\ rad}$. This is significantly smaller than our original target of $10^{-1}$ rad, and should be sufficiently accurate for the analysis of EMRI signals that LISA is expected to observe. We emphasise that this extrapolation assumes that our interpolation errors are significantly larger than the numerical precision errors in our computed data. In general, the worst-behaved systems have very high spins ($a>0.99$) and/or very high initial eccentricities ($e_{0}>0.8$). This highlights that these systems may require a more careful treatment (both in the generation and interpolation of forcing function data), an exploration of which we leave for future work.
It is also useful to understand whether insufficient sampling density in one of $u$, $w$ or $z$ (corresponding to $p$, $e$ and $a$) contributes significantly to interpolation error in $\hat{f}_{p,e}^{(0)}$. To examine this more closely, we also construct trajectory models from grids down-sampled by a factor of $2$ in a single dimension at at time. Histograms of dephasings between trajectories from these models and our full-resolution model are shown in the bottom panel of Fig. 8. As these histograms overlap significantly, we conclude that the point density of our grids is well-proportioned with respect to grid dimension. As expected, these histograms also lie between those corresponding to the extrapolated dephasings and the model constructed with grids down-sampled by $2$ in all dimensions.

### v2

@@ III.3 Mode selection and waveform summation
While the inclusion of spin on the primary introduces additional complexity to the generation of inspiral trajectories and mode amplitudes, once these have been addressed, the waveform may be constructed in much the same way as for spinless systems. We review this procedure below, which is largely unchanged except for the incorporation of the continuous ODE solution described in Section˜III.1.3; see Katz:2021yft and Speri:2023jte for more information regarding these stages of waveform generation in the time domain and frequency domain, respectively.
@@
Before performing the summation in Eq.˜1, we apply a mode selection process to reduce the number of modes that must be summed without sacrificing accuracy, which proceeds as follows. First, at each sparse trajectory point, we compute the mode power $|A_{\ell mn}|^{2}$ for each waveform mode in our amplitude data grids. We then sort the list of mode power values at each individual point, perform a cumulative summation over this list and truncate the sum once it exceeds a fraction $(1-\kappa)$ of the total power. We assume that modes that do not pass this cut-off do not strongly contribute to the overall waveform. However, to ensure that the waveform is smooth with respect to time, we take the union of the sets of modes that pass the individual cut-offs and retain only these modes for waveform summation. The tuning parameter $\kappa$ controls the accuracy of the waveform with respect to this procedure; in line with previous work, we set $\kappa=10^{-5}$ by default. We have verified that this value is a sensible choice for a variety of EMRIs and IMRIs, as discussed in Section˜V.6.
@@
After mode selection, we perform the summation in Eq.˜1 over the remaining sets of mode indices. For time-domain waveforms, we interpolate the sparse quantities $\mathcal{A}_{\ell mn}$ and $\Phi_{mn}$ (with cubic splines and the continuous ODE solution respectively) on a uniform grid of time values of spacing $\mathrm{d}t$, summing the interpolated waveform modes together according to Eq.˜1. Frequency-domain waveforms are obtained via the approach described in Speri:2023jte , making use of the stationary phase approximation. As this approach assumes that the frequency evolution of each mode is described by cubic splines (such that the stationary phase relation has analytic roots), we retain this structure rather than incorporating the continuous solution depicted in Fig.˜2. The frequency-domain waveform also requires higher-order frequency derivatives to approximate the Fourier phase in each bin; as the derivatives of this cubic spline are less accurate than those obtained with the continuous solution, we rewrite this part of the frequency-domain summation to use the coefficients output by the ODE solver. We found that this significantly improved the consistency of the time-domain and frequency-domain waveforms. Modifying the root-finding procedure of the stationary phase approximation to incorporate our more accurate solution for the frequency evolution (e.g. by taking the analytic root as an input to Newton’s method, which will converge in 1-2 iterations) will further improve this consistency at minimal computational cost, and is left to future work.
@@ IV Validation of model accuracy
With our waveform generation framework fully specified, we will now explore the accuracy of our waveform model at adiabatic order. This is made difficult by lack of accurate waveform templates to compare against: numerical relativity simulations have not been computed (however, see recent efforts in Refs. Lousto:2022hoq ; Wittek:2024gxn ; Wittek:2024pis ) in this region of the parameter space (especially for the number of orbital cycles $\sim\mathcal{O}(\nu^{-1})$] required for useful comparisons to be made), and PN models are not trustworthy in the strong-field region of our domain where confirming the accuracy of our model is most pertinent. We therefore opt for a different approach in which we quantify significant sources of systematic error in our model (which in combination are responsible for any differences between our model and “exact” adiabatic EMRI waveforms), with the aim of understanding how these errors propagate to the accuracy of our constructed waveforms. In this way, we can roughly quantify the faithfulness of our model with respect to an “error-free” adiabatic model without the need for independent waveforms to compare against. Throughout this section, we will examine each component of our waveform generation in turn: we first consider the trajectory module (Section˜IV.1), followed by the amplitude module (Section˜IV.2) and conclude with an examination of how these systematics impact accuracy at the waveform template level (Section˜IV.3).
In addition to this self-contained examination of our model, we also verify that it correctly matches other models in limiting cases in Appendix˜E. In combination with the systematic validation that we perform throughout this section, the results of these comparisons confirm that our model produces accurate waveforms for spinning and/or equatorial IMRIs and EMRIs across its domain of validity of $|a|\leq 0.999$ and $e_{0}\leq 0.9$. In this process, we identify that our model is least accurate for: $a>0.99$ and $p-p_{\mathrm{sep}}\lesssim 1$; and $e\gtrsim 0.85$. These are regions where self-force data vary rapidly and are most challenging to interpolate.
@@ IV.1 Trajectory validation
In this subsection, we verify that our model for the inspiral trajectory of the CO accurately traces the orbital phasing of the system to adiabatic order over the long timescales typical of the asymmetric-mass binary systems observable by LISA. As our model only incorporates adiabatic contributions to the inspiral evolution, one could argue that we need only construct the phase evolution to an accuracy of $\mathcal{O}(1)$ radian, as we cannot hope to match trajectories that include post-adiabatic contributions to better than this scale (and typically, ignoring post-adiabatic terms regularly leads to 1010-$100$ radians of phase error Isoyama:2021jjd ; Wardell:2021fyy ; Burke:2023lno ; Lynch:2024ohd , at least for the quasi-circular EMRIs). However, we can assume that post-adiabatic effects will be implemented in our framework once they are readily obtainable (as they are necessary to achieve sufficient accuracy for data analysis purposes), so this is not a useful requirement for us to satisfy. In this work, our focus remains on the accuracy of our model for the adiabatic contribution to the inspiral. Quantitatively assessing how phase errors translate into systematic biases in EMRI parameter estimation is not straightforward. However, in Ref. Burke:2023lno , which examined quasi-circular inspirals into non-spinning MBHs, a phase error of $\sim 10^{-1}$ radians was sufficiently small for any resulting biases to remain subdominant to statistical uncertainties. Our aim therefore is to ensure that our trajectory model reproduces inspiral phasing with at least this level of accuracy.
@@
Figure 5: Relative error in $\hat{f}_{p}^{(0)}$ between an interpolation of our data grid and the corresponding data of Ref. Fujita:2020zxe , available from the BHPC data lake BHPClub . The two datasets agree to better than 88 digits over the majority of the parameter space. Here we examine prograde orbits with $a=0.5$; we observe similar agreement for other spin values $a\in[-0.9,0.9]$.
@@
The first of these datasets was computed by the authors of Ref. Fujita:2020zxe (built on the BHPC code developed in Refs. Fujita:2004rb ; Fujita:2005kng ; Fujita:2009us ) and is publicly available on the BHPC website BHPClub . It broadly and densely covers a large fraction of our parameter space. The authors accompany their data with an error estimate informed by their truncation of mode-summations. For $p\lesssim p_{\mathrm{sep}}+1$, the estimated error becomes prohibitively large for validation purposes — in this region, the discrepancy between our data and theirs grows proportionally to the predicted error. We therefore excise this region of the parameter space from this comparison (and will examine it separately below with a separate dataset). The region of the BHPC data we retain extends to $p\sim 25$ and $e=0.925$, providing good coverage of a large fraction of our grid. Beyond $p\sim 25$, we validate our data separately with BHPC’s PN results (see Section˜E.2). The relative error between an interpolation of our data and the BHPC data for $\hat{f}_{p}^{(0)}$ is shown in Fig.˜5 — for brevity, we do not include the comparison with $\hat{f}_{e}^{(0)}$, which follows similar behaviour. We attain excellent agreement with this dataset across the parameter space, with relative error remaining below $\sim 10^{-8}$ over the vast majority of the grid. Above $e\sim 0.8$, errors steadily rise to $\sim 10^{-4}$. This is expected: mode-sums with respect to $n$ converge slowly in this region, requiring one to compute and sum over hundreds of $n$-modes. The large $n$-harmonics become increasingly more difficult to resolve at floating point precision and introduce their own source of error if they are under-resolved (similar to how it is incredibly difficult to calculate the 100th harmonic in a Fourier series expansion). For a single $n$-mode, this error is typically subdominant to the total error, but if the mode-sum is not truncated early enough, the accumulation of these small errors can saturate the overall relative accuracy of the results. As both our data and that of the BHPC are impacted by this limitation, it is not clear which of the two have been computed more accurately in this high eccentricity region, only that the two datasets are consistent to $\sim 4$ digits of precision. This comparison is performed for prograde orbits with $a=0.5$ — we find similarly good agreement across all spin values $(a=0.1,0.3,0.5,0.7,0.9)$ for both prograde and retrograde configurations, confirming that we have accurately computed and interpolated our $\hat{f}_{p,e}^{(0)}$ data grids over this part of the parameter space.
[…]
@@
To inspect the accuracy of our data grids in the strong-field region $p\rightarrow p_{\mathrm{sep}}$, we also compare against data computed with the GREMLIN code, independently developed by author Hughes and previous collaborators. (See Ref. Drasco:2005kz for an overview of this code’s foundations, and Refs. ThroweThesis ; OSullivan:2014ywd for discussion of updates to notation and methods that this code uses. Additional recent optimizations and algorithmic improvements, including a port to the Open Science Grid osg07 , will be described in a future publication, accompanied by a public release after removing proprietary code currently used by this package.) This data is more densely sampled in this region than our grids in $a$ and $p$; comparisons are done on $4$ slices in eccentricity that follow a taper similar to the one we include in our grid reparametrization. We distinguish forcing functions computed with few and GREMLIN with superscript “FEW” and “SAH” respectively. The relative error between an interpolation of our data and each of these slices is shown in Fig.˜6 for the case of $\hat{f}_{p}^{(0)}$. We observe excellent agreement (with relative errors less than one part in a million) for $p-p_{\mathrm{sep}}\gtrsim 10^{-2}$. Closer to the separatrix, we observe a larger discrepancy between the datasets, with relative errors reaching $\sim 10^{-3}$; this is expected as (especially at larger eccentricities) it is computationally challenging to compute $\hat{f}_{p,e}^{(0)}$ in this region, and our interpolation error begins to dominate at high values of $a$ (as indicated by the alternating bands of larger error with respect to $a$), as we will examine further below. However, EMRIs typically complete relatively few orbital cycles in this region of the parameter space (compared to their total number of in-band cycles) before crossing the separatrix, so even relatively large errors will not significantly impact the phase accuracy of the trajectory.
@@
While the deviations observed are due in part to errors in the interpolation of our data grids, they are also impacted by differences in the convergence criteria associated with each software. The gremlin computations truncate at either a relative precision of $10^{-7}$ or up to $n=250$ closest to the separatrix, whereas pybhpt cuts off at either a relative precision of $10^{-8}$ or $n=300$; as these higher $n$-modes can contain significant errors due to limited numerical precision, truncating the mode-sum at different values of $n$ may yield results that differ substantially beyond the first few digits. Moreover, relative precision cut-offs may not be directly comparable between datasets, as this quantity may be estimated in different numerical frameworks. Likewise, the BHPC data were computed with mode-sums truncated at up to $n=1000$ or $\ell=25$ for the (highly-eccentric) strongest-field orbits to ensure a relative precision at a level of $10^{-6}$ Fujita:2020zxe , which may again yield significant differences with our data (such as those observed for the highest eccentricities in Fig.˜5). Errors due to these differences in methodology as a function of parameter space are not well understood at present; further investigation of these systematics is beyond the scope of this work and reported in a separate work Khalvati:2025znb (but is worthy of future study, particularly for eccentric and inclined orbits).
@@
Two other significant features are present in the heatmaps shown in Fig.˜6. The first is that for $a\rightarrow 0.999$, we observe a significant increase in relative error in alternating bands of constant $a$. These bands correspond to $a$-values lying on grid nodes (low interpolation error) and exactly between grid nodes (high interpolation error) respectively, and their existence suggests that our grid density with respect to $a$ is insufficient near $p_{\mathrm{sep}}$ (even despite our reparameterization, which concentrates grid points in this region). As we will see later in this subsection, this has little impact on the accuracy of our trajectories. The second feature is observed for the largest eccentricity slice (bottom-right panel of Fig.˜6), $a\in[0,0.5]$ and $p-p_{\mathrm{sep}}\lesssim 1$, where a patch of larger relative error is observed. These points correspond to $e\sim 0.5$ near the separatrix (due to the eccentricity taper with respect to $e_{\mathrm{max}}=0.9$, see Appendix˜B) and the larger errors are a consequence of ill-converged mode-sums typical of more eccentric orbits near $p=p_{\mathrm{sep}}$. As this eccentricity slice is at the edge of our grid, trajectories typically do not perform many orbital cycles in this region before their eccentricities decay to lower values where this numerical precision issue is absent; we therefore do not anticipate this will significant impact the accuracy of our model. In this eccentricity slice, a steadily-growing relative error is also observed for $a\rightarrow-0.999$, which corresponds to $e\rightarrow 0.9$ and is explained by the differences in convergence criteria between implementations (as described above).
@@
Taking these comparisons with independent datasets as a reference point, we assess the impact of interpolation error on our estimation of $f_{p,e}^{(0)}$ as a function of the inspiral parameters. As our data grids have dimensions of the form $2^{n}+1$ for some positive integer $n$, we can construct a grid of dimension $2^{n-1}+1$ (halving the point density) by omitting every second point in the grid, which we will refer to as “down-sampling” the grid by a factor of $2$. We can then evaluate this down-sampled grid at the omitted points, placing the upper bound on the interpolation error for the full grid across the parameter space. The comparison is performed in Fig.˜7 for $a\sim 0.998$, the spin value at which we find the interpolation errors to be most significant (in line with what is observed in Fig.˜6); again, we omit our comparison with $\hat{f}_{e}^{(0)}$ for brevity, which yields similar error estimates. Over the majority of the parameter space, this interpolation error estimate exceeds those shown in Figs.˜6 and 5 where we compare to independent datasets. This means that for a trajectory model built with data grids down-sampled by a factor of 2 (or higher), interpolation error due to insufficient point density will be the dominant contribution to orbital dephasing.
@@
We will now use this result to roughly quantify the accuracy of our model. We consider models built from data-grids down-sampled by factors of $2$, 44 and $8$, and compare phase trajectories constructed with these down-sampled models to those produced by our full-scale model (which has the full grid density). In the top panel of Fig.˜8, we plot histograms for dephasings between each down-sampling scenario and our full-resolution model for a random sample of $10^{3}$ inspirals with $p_{0}$ tuned such that they inspiral after 4 years (see Appendix˜C for details of the random-draw procedure). We observe that these dephasings display power-law convergence (i.e., in line with the power of $2$ of the down-sampling factor) as a function of grid density. By fitting this convergence trend between grid density and dephasing for each data point, we then extrapolate to obtain an estimate of the phase error at the full grid density, resulting in the dashed histogram in the top panel of Fig.˜8. Taking this extrapolation as a rough estimate of our model’s dephasing with respect to a “perfect” adiabatic inspiral model, we predict that our model attains a phase accuracy of $\sim 10^{-4}\mathrm{\penalty 10000\ rad}$ over the majority of the parameter space, with a higher tail extending to $\sim 10^{-2}\mathrm{\penalty 10000\ rad}$. This is significantly smaller than our original target of $10^{-1}$ rad, and should be sufficiently accurate for the analysis of EMRI signals that LISA is expected to observe. We emphasise that this extrapolation assumes that our interpolation errors are significantly larger than the numerical precision errors in our computed data. In general, the worst-behaved systems have very high spins ($a>0.99$) and/or very high initial eccentricities ($e_{0}>0.8$). This highlights that these systems may require a more careful treatment (both in the generation and interpolation of forcing function data), an exploration of which we leave for future work.
It is also useful to understand whether insufficient sampling density in one of $u$, $w$ or $z$ (corresponding to $p$, $e$ and $a$) contributes significantly to interpolation error in $\hat{f}_{p,e}^{(0)}$. To examine this more closely, we also construct trajectory models from grids down-sampled by a factor of $2$ in a single dimension at at time. Histograms of dephasings between trajectories from these models and our full-resolution model are shown in the bottom panel of Fig.˜8. As these histograms overlap significantly, we conclude that the point density of our grids is well-proportioned with respect to grid dimension. As expected, these histograms also lie between those corresponding to the extrapolated dephasings and the model constructed with grids down-sampled by $2$ in all dimensions.